# From Many Models to ONE THEORY From the Dark Ages to Enlightenment J. Iliopoulos

### v1

@@ 2 Elementary Particles: The origins
Λετ υς ςταρτ ωιτη ΧΕΔ. Τηε ιναριανςε υνδερ γαυγε τρανςφορματιονς ιμπλιες τηατ, ιφ ωε φινδ α ςολυτιον φορ τηε εςτορ ποτεντιαλ $A_{\mu}(x)=\hat{A}_{\mu}(x)$, ωε ςαν ςονςτρυςτ αν ινφινιτψ οφ οτηερς βψ ωριτινγ $A_{\mu}(x)=\hat{A}_{\mu}(x)+\partial_{\mu}\theta(x)$ ωιτη $\theta(x)$ αν αρβιτραρψ φυνςτιον. Τηε γαυγε ςονδιτιονς αρε ςυπποςεδ το ρεςτριςτ τηις αρβιτραρινεςς ανδ, ηοπεφυλλψ, ςηοοςε α υνιχυε ςολυτιον [^26] V.N. Gribov has shown that, in general, there exists no gauge condition yielding a unique solution[41]. This phenomenon is known as “Gribov ambiguity”. It is important in the attempts to give a global definition of a gauge theory, but it does not affect the perturbation expansion.. Σινςε ωε ϰνοω τηατ ιν Νατυρε ονλψ τρανςερςε πηοτονς αρε πηψςιςαλ, τηε οβιους ιδεα ις το ιμποςε τηε ῝ουλομβ γαυγε ςονδιτιον ${\bm{\nabla}}\cdot{\bm{A}}(x)\!=\!0$. Ιτ ις εαςψ το ςηοω τηατ ιν τηις γαυγε ωε αρε λεφτ ωιτη τηε τωο “πηψςιςαλ” δεγρεες οφ φρεεδομ, το ωιτ τηε τρανςερςε πηοτονς. Ιτ ις τηε μετηοδ υςεδ βψ Διρας ιν τηε ςαλςυλατιον οφ τηε ςποντανεους εμιςςιον προβαβιλιτψ, τηε φιρςτ πηψςιςαλ αππλιςατιον οφ τηε χυαντιςεδ ραδιατιον φιελδ[42].
@@
ϒςινγ μοδερν νοτατιον, ωε ςαν ωριτε τηε ρενορμαλιςατιον γρουπ εχυατιον φορ τηε ςιμπλεςτ ςςαλαρ φιελδ τηεορψ οφ ρεφερενςε (11) ας φολλοως: Τηε πηψςιςαλ τηεορψ, ι.ε. τηε ονε ωηιςη ςαν βε υςεδ το ςομπαρε ωιτη τηε εξπεριμενταλ δατα, ις τηε ονε ιν ωηιςη τηε αλυε οφ τηε μαςς εχυαλς τηε πηψςιςαλλψ μεαςυρεδ αλυε $m_{\mathrm{ph}}$, τηε φιελδ ις νορμαλιςεδ το ονε ον τηε μαςς-ςηελλ, ανδ τηε ςουπλινγ ςονςταντ ις τηε αλυε οφ τηε φουρ-ποιντ φυνςτιον ατ ςομε πηψςιςαλλψ μεαςυραβλε ποιντ, φορ εξαμπλε τηε ςςαττερινγ αμπλιτυδε ατ τηρεςηολδ. Τηε ςορρεςπονδινγ τηρεε ρενορμαλιςατιον ςονδιτιονς ταϰε τηε φορμ [^36] The renormalisation program is usually formulated in terms of the “amputated, one-particle-irreducible $n$-particle correlation functions” $\Gamma^{(n)}$. They are defined as the sum of all connected diagrams with $n$ external lines which cannot be separated into two disconnected parts by cutting a single internal line (the meaning of “one-particle-irreducible”) and from which we have dropped the propagators of the external lines (the meaning of “amputated”).
@@
Τηε ςονςεπτ οφ τηε ρενορμαλιςατιον γρουπ ςτεμς φρομ τηε ρεαλιςατιον τηατ, φρομ τηε πυρελψ τεςηνιςαλ ποιντ οφ ιεω, νοτηινγ φορςες υς το ςηοοςε τηεςε πηψςιςαλ ρενορμαλιςατιον ςονδιτιονς· ανψ οτηερ ςηοιςε ωουλδ βε αςςεπταβλε, προιδεδ ιτ ψιελδς ωελλ-δεφινεδ αλυες φορ τηε τηρεε αρβιτραρψ ςονςταντς. Τηε εξαμπλε ωηιςη ις μοςτ ςομμονλψ υςεδ ις το ιντροδυςε α μαςς παραμετερ $M$ ανδ ρεωριτε τηε ςεςονδ ανδ τηιρδ οφ τηε εχυατιονς (29 ας [^37] We could use even more general conditions, for example not keeping the physical mass in the first of the equations (29), but no substantially new insight is gained. Examples of such more general renormalisation group equations are given in reference [72].
@@
Τηε πηψςιςαλ μεανινγ οφ (36) ις ςλεαρ. Σςαλινγ αλλ μομεντα οφ α Γρεεν φυνςτιον βψ α ςομμον φαςτορ $\rho$ ανδ ταϰινγ $\rho$ λαργε, ηας τηε φολλοωινγ εφφεςτς: (ι) Ιτ μυλτιπλιες εερψ εξτερναλ λινε βψ τηε ςςαλε δεπενδεντ φαςτορ $\exp(-\int_{1}^{\rho}{\rm d}\rho^{\prime}\gamma(\bar{\lambda}(\lambda,\rho^{\prime})))$414141 We see that the Green functions get rescaled as if they had dimensions related to $\gamma(\lambda)$. This is the origin of the term “anomalous dimension”. ανδ (ιι) ιτ ρεπλαςες τηε πηψςιςαλ ςουπλινγ ςονςταντ $\lambda$ βψ αν εφφεςτιε ονε, $\bar{\lambda}=\bar{\lambda}(\lambda,\rho)$, ωηιςη ις τηε ςολυτιον οφ (35). Φορ τηις ρεαςον $\bar{\lambda}$ ις οφτεν ςαλλεδ ρυννινγ ςουπλινγ ςονςταντ.
@@
Ιν ςεςτιον 1 ωε πρεςεντεδ τηε ςονςτρυςτιον οφ τηε ιςοςπιν ςψμμετρψ οφ νυςλεαρ φορςες βψ Ηειςενβεργ ανδ Κεμμερ. Ιτ ις βαςεδ ον τηε γρουπ $SU(2)$ ανδ τηε παιρ προτον-νευτρον βελονγς το τηε τωο-διμενςιοναλ φυνδαμενταλ ρεπρεςεντατιον. Τηε διςςοερψ οφ ςτρανγε παρτιςλες βρουγητ τηε νεεδ το ενλαργε τηις ςψμμετρψ ανδ τηε “οβιους” εξτενςιον ωας $SU(2)\rightarrow SU(3)$ βψ ενλαργινγ τηε δουβλετ το α τριπλετ. Ιτ ωας τηε μοδελ προποςεδ βψ Σ. Σαϰατα ιν 1956[77] ιν ωηιςη ηε αδδεδ τηε $\Lambda$-ηψπερον το τηε τωο νυςλεονς το φορμ α φυνδαμενταλ τριπλετ οφ $SU(3)$. Τηις ςςηεμε διδ νοτ φιτ τηε οβςερεδ ηαδρονς ανδ ιν τηε φολλοωινγ ψεαρς μανψ γρουπς ωερε ςονςιδερεδ ας ςανδιδατες φορ τηις λαργερ ςψμμετρψ. Φιναλλψ ιτ τυρνεδ ουτ τηατ Νατυρε ωας μορε ςυβτλε: $SU(3)$ ις τηε ριγητ ανςωερ, βυτ τηε φυνδαμενταλ τριπλετ ις νοτ φορμεδ βψ παρτιςλες ωε ςεε. Ιν 1961 Μ. Γελλ-Μανν ανδ Ψ. Νεέμαν φορμυλατεδ τηε μοδελ ωηιςη βεςαμε ϰνοων ας “Τηε ειγητφολδ ωαψ”[78] ιν ωηιςη τηε λοω-λψινγ ηαδρονς φορμ οςτετς οφ $SU(3)$. Τηε πεςυλιαρ προπερτψ οφ τηις ςςηεμε ις πρεςιςελψ τηε αβςενςε οφ πηψςιςαλ παρτιςλες βελονγινγ το τηε φυνδαμενταλ τριπλετ ρεπρεςεντατιον. Τηις λεδ Γελλ-Μανν ανδ, ινδεπενδεντλψ Γ. Ζωειγ, το προποςε α ςομποςιτε μοδελ ιν ωηιςη ηαδρονς αρε βουνδ ςτατες οφ ςονςτιτυεντς – ναμεδ “χυαρϰς” βψ Γελλ-Μανν ανδ “αςες” βψ Ζωειγ – ωηιςη φιλλ τηε μιςςινγ τριπλετ[79]. Ωηατ ις λεςς ϰνοων ις τηατ ςιμιλαρ ιδεας ηαδ βεεν εξπρεςςεδ εαρλιερ, Ι βελιεε φιρςτ βψ Λ.Β. Οϰυν[80]. Ηε ςταρτεδ φρομ τηε Σαϰατα μοδελ βυτ ηε ιντερπρετεδ τηε μεμβερς οφ τηε φυνδαμενταλ τριπλετ νοτ ας τηε πηψςιςαλ $p$, $n$ ανδ $\Lambda$, βυτ ας φιςτιτιους παρτιςλες [^44] He used the term “sakatons”.. Ιν φαςτ, ουρ ιδεας ον χυαρϰς φολλοωεδ α ςιμιλαρ τραϑεςτορψ. Ατ φιρςτ τηεψ ωερε ςονςιδερεδ ας ϑυςτ ανοτηερ λαψερ οφ ςομποςιτενεςς ανδ εξπεριμεντς ωερε δεςιγνεδ το διςςοερ τηεμ [^45] There have been even several unsubstantiated claims of discovery.. Ονε οφ τηε μοτιατιονς το βυιλτ ΙΣΡ, τηε φιρςτ ηαδρονις ςολλιδερ, ωας το διςςοερ χυαρϰς. Ηοωεερ, τηε πιςτυρε βεςαμε ςοον μορε ςομπλιςατεδ. Τηε νεεδ το ςατιςφψ τηε ρεχυιρεμεντς οφ Φερμι ςτατιςτιςς, λεδ το τηε ιντροδυςτιον οφ ςολορ[81] ανδ τηε ηψποτηεςις οφ “ςολορ βλινδνεςς” λεδ το τηε νοτιον οφ ςονφινεμεντ. Εεν τοδαψ ωε ηαε νο ρεαλ υνδερςτανδινγ οφ ιτ ανδ ωε ςαννοτ ριγορουςλψ δεριε ιτ φρομ φιρςτ δψναμιςαλ πρινςιπλες [^46] It is one of the unsolved problems endowed with a prize of one million dollars from the Clay Mathematics Institute. For a review of the various ideas around confinement see reference [82].. Ας α ρεςυλτ, ιν ςπιτε οφ ςομε πηενομενολογιςαλ ςυςςεςςες οφ τηε ςιμπλε νον-ρελατιιςτις χυαρϰ μοδελ, μοςτ πεοπλε ςηιφτεδ τηειρ ιεως το ςονςιδερινγ χυαρϰς ας μερε ματηεματιςαλ δειςες το ϰεεπ τραςϰ οφ ςψμμετρψ προπερτιες. Νατυραλλψ, Μ. Γελλ-Μανν, ωηο ςτρονγλψ αδοςατεδ τηις οπινιον, πλαψεδ α μαϑορ ρολε το ιτς γενεραλ ενδορςεμεντ. Τηε ατμοςπηερε ςηανγεδ ωιτη τηε ΣΛΑ῝ ρεςυλτς ον δεεπ ινελαςτις ςςαττερινγ, βυτ Ι ωιλλ διςςυςς τηεμ ιν ςεςτιον 1.
@@
Ι βελιεε τηατ τηε φιρςτ περςον ωηο δεςιδεδ το λοοϰ ατ τηε Ψανγ-Μιλλς χυαντυμ φιελδ τηεορψ ας τηε δψναμιςαλ τηεορψ οφ ωεαϰ ιντεραςτιονς ωας Μ.Θ.Γ. ῞ελτμαν. Αρουνδ 1966, ηε ωας τρψινγ το υνδερςτανδ τηε δεεπερ οριγιν οφ τηε ςονςερατιον, ορ νεαρ ςονςερατιον, οφ τηε ωεαϰ ςυρρεντς. Ιν παρτιςυλαρ, ηε τριεδ το τηροω ςομε λιγητ ον τηε γενεραλ ςονφυςιον ωηιςη πρεαιλεδ ατ τηατ τιμε ςονςερνινγ τηε ςο-ςαλλεδ “Σςηωινγερ τερμς” ιν τηε ςομμυτατορς οφ τωο ςυρρεντ ςομπονεντς. Ωηιλε ηε ωας ον α ιςιτ φρομ ῝ΕΡΝ το Βροοϰηαεν, ηε ωροτε α παπερ ιν ωηιςη ηε ςυγγεςτεδ α ςετ οφ διεργενςε εχυατιονς ωηιςη γενεραλιςεδ τηε νοτιον οφ τηε ςοαριαντ δεριατιε οφ χυαντυμ ελεςτροδψναμιςς. Τηις φυνδαμενταλ ιδεα ωας ταϰεν υπ νεξτ ψεαρ ανδ δεελοπεδ φυρτηερ βψ Θ.Σ. Βελλ[117]. Τηεςε εχυατιονς λοοϰεδ λιϰε τηε ςοαριαντ δεριατιες οφ νον-αβελιαν γαυγε τηεοριες, ςο ῞ελτμαν δεςιδεδ το ςτυδψ τηειρ φιελδ τηεορψ προπερτιες. Τηε ελεςτροδψναμιςς οφ ςηαργεδ εςτορ βοςονς ηαδ βεεν φορμυλατεδ αλρεαδψ βψ Τ.Δ. Λεε ανδ ῝.Ν. Ψανγ ιν 1962[118]. Τηεψ ηαδ ςηοων τηατ ελεςτρομαγνετις γαυγε ιναριανςε αλλοως το εξπρεςς τηε εςτορ βοςον´ς ςηαργε $e$, μαγνετις μομεντ $\mu$ ανδ χυαδρυπολε μομεντ $Q$ ιν τερμς οφ ονλψ τωο παραμετερς $e$ ανδ $\kappa$, ας $\mu=e(1+\kappa)/2m_{W}$ ανδ $Q=-e\kappa/m_{W}^{2}$. Τηε ρεςυλτινγ τηεορψ ις ηιγηλψ διεργεντ βυτ ῞ελτμαν νοτιςεδ τηατ μανψ διεργενςες ςανςελ φορ τηε παρτιςυλαρ αλυε $\kappa=1$. Ιτ ις τηε αλυε πρεδιςτεδ βψ α τηεορψ ιν ωηιςη $W^{\pm}$ ανδ τηε πηοτον φορμ α Ψανγ-Μιλλς τριπλετ. Φορ ῞ελτμαν τηις ωας α ςλεαρ ςιγναλ τηατ τηε τηεορψ οφ ωεαϰ ανδ ελεςτρομαγνετις ιντεραςτιονς μυςτ οβεψ α Ψανγ-Μιλλς γαυγε ιναριανςε.
Στυδψινγ τηε ρενορμαλιςατιον προπερτιες οφ Ψανγ-Μιλλς τηεοριες τυρνεδ ουτ το βε α ερψ ηαρδ ανδ ςομπλεξ προβλεμ, βοτη ςονςεπτυαλλψ ανδ πραςτιςαλλψ. Ον τηε ςονςεπτυαλ ςιδε, νειτηερ τηε ρεςυλτς οφ Φεψνμαν[52] ανδ Φαδδεε ανδ Ποπο[53] νορ τηοςε ον ςποντανεους ςψμμετρψ βρεαϰινγ ωερε ωιδελψ ϰνοων ανδ ῞ελτμαν ηαδ το ρεδιςςοερ εεν τηε βαςις Φεψνμαν ρυλες. Ον τηε πραςτιςαλ ςιδε, τηε νυμβερ οφ τερμς γρεω ερψ φαςτ ανδ ῞ελτμαν ηαδ το δεελοπ α ςομπυτερ προγραμ το ηανδλε τηεμ[119]. Ηε ςαλλεδ ιτ “Σςηοονςςηιπ”,606060“Clean ship” in dutch ανδ ιτ ωας τηε φιρςτ προγραμ οφ ςψμβολις μανιπυλατιονς αππλιεδ το τηεορετιςαλ ηιγη ενεργψ πηψςιςς.

### v2

@@ 2 Elementary Particles: The origins
Λετ υς ςταρτ ωιτη ΧΕΔ. Τηε ιναριανςε υνδερ γαυγε τρανςφορματιονς ιμπλιες τηατ, ιφ ωε φινδ α ςολυτιον φορ τηε εςτορ ποτεντιαλ $A_{\mu}(x)=\hat{A}_{\mu}(x)$, ωε ςαν ςονςτρυςτ αν ινφινιτψ οφ οτηερς βψ ωριτινγ $A_{\mu}(x)=\hat{A}_{\mu}(x)+\partial_{\mu}\theta(x)$ ωιτη $\theta(x)$ αν αρβιτραρψ φυνςτιον. Τηε γαυγε ςονδιτιονς αρε ςυπποςεδ το ρεςτριςτ τηις αρβιτραρινεςς ανδ, ηοπεφυλλψ, ςηοοςε α υνιχυε ςολυτιον [^27] V.N. Gribov has shown that, for Yang-Mills theories, there exists no gauge condition yielding a unique solution[41]. This phenomenon is known as “Gribov ambiguity”. It is important in the attempts to give a global definition of a gauge theory, but it does not affect the perturbation expansion.. Σινςε ωε ϰνοω τηατ ιν Νατυρε ονλψ τρανςερςε πηοτονς αρε πηψςιςαλ, τηε οβιους ιδεα ις το ιμποςε τηε ῝ουλομβ γαυγε ςονδιτιον ${\bm{\nabla}}\cdot{\bm{A}}(x)\!=\!0$. Ιτ ις εαςψ το ςηοω τηατ ιν τηις γαυγε ωε αρε λεφτ ωιτη τηε τωο “πηψςιςαλ” δεγρεες οφ φρεεδομ, το ωιτ τηε τρανςερςε πηοτονς. Ιτ ις τηε μετηοδ υςεδ βψ Διρας ιν τηε ςαλςυλατιον οφ τηε ςποντανεους εμιςςιον προβαβιλιτψ, τηε φιρςτ πηψςιςαλ αππλιςατιον οφ τηε χυαντιςεδ ραδιατιον φιελδ[42].
@@
ϒςινγ μοδερν νοτατιον, ωε ςαν ωριτε τηε ρενορμαλιςατιον γρουπ εχυατιον φορ τηε ςιμπλεςτ ςςαλαρ φιελδ τηεορψ δεςςριβεδ βψ τηε Λαγρανγιαν δενςιτψ (11) ας φολλοως: Τηε πηψςιςαλ τηεορψ, ι.ε. τηε ονε ωηιςη ςαν βε υςεδ το ςομπαρε ωιτη τηε εξπεριμενταλ δατα, ις τηε ονε ιν ωηιςη τηε αλυε οφ τηε μαςς εχυαλς τηε πηψςιςαλλψ μεαςυρεδ αλυε $m_{\mathrm{ph}}$, τηε φιελδ ις νορμαλιςεδ το ονε ον τηε μαςς-ςηελλ, ανδ τηε ςουπλινγ ςονςταντ ις τηε αλυε οφ τηε φουρ-ποιντ φυνςτιον ατ ςομε πηψςιςαλλψ μεαςυραβλε ποιντ, φορ εξαμπλε τηε ςςαττερινγ αμπλιτυδε ατ τηρεςηολδ. Τηε ςορρεςπονδινγ τηρεε ρενορμαλιςατιον ςονδιτιονς ταϰε τηε φορμ [^37] The renormalisation program is usually formulated in terms of the “amputated, one-particle-irreducible $n$-particle correlation functions” $\Gamma^{(n)}$. They are defined as the sum of all connected diagrams with $n$ external lines which cannot be separated into two disconnected parts by cutting a single internal line (the meaning of “one-particle-irreducible”) and from which we have dropped the propagators of the external lines (the meaning of “amputated”).
@@
Τηε ςονςεπτ οφ τηε ρενορμαλιςατιον γρουπ ςτεμς φρομ τηε ρεαλιςατιον τηατ, φρομ τηε πυρελψ τεςηνιςαλ ποιντ οφ ιεω, νοτηινγ φορςες υς το ςηοοςε τηεςε πηψςιςαλ ρενορμαλιςατιον ςονδιτιονς· ανψ οτηερ ςηοιςε ωουλδ βε αςςεπταβλε, προιδεδ ιτ ψιελδς ωελλ-δεφινεδ αλυες φορ τηε τηρεε αρβιτραρψ ςονςταντς. Τηε εξαμπλε ωηιςη ις μοςτ ςομμονλψ υςεδ ις το ιντροδυςε α μαςς παραμετερ $M$ ανδ ρεωριτε τηε ςεςονδ ανδ τηιρδ οφ τηε εχυατιονς (29) ας [^38] We could use even more general conditions, for example not keeping the physical mass in the first of the equations (29), but no substantially new insight is gained. Examples of such more general renormalisation group equations are given in reference [72].
@@
Τηε πηψςιςαλ μεανινγ οφ (36) ις ςλεαρ. Σςαλινγ αλλ μομεντα οφ α Γρεεν φυνςτιον βψ α ςομμον φαςτορ $\rho$ ανδ ταϰινγ $\rho$ λαργε, ηας τηε φολλοωινγ εφφεςτς: (ι) Ιτ μυλτιπλιες εερψ εξτερναλ λινε βψ τηε ςςαλε δεπενδεντ φαςτορ $\exp(-\int_{1}^{\rho}{\rm d}\rho^{\prime}\gamma(\bar{\lambda}(\lambda,\rho^{\prime})))$424242 We see that the Green functions get rescaled as if they had dimensions related to $\gamma(\lambda)$. This is the origin of the term “anomalous dimension”. ανδ (ιι) ιτ ρεπλαςες τηε πηψςιςαλ ςουπλινγ ςονςταντ $\lambda$ βψ αν εφφεςτιε ονε, $\bar{\lambda}=\bar{\lambda}(\lambda,\rho)$, ωηιςη ις τηε ςολυτιον οφ (35). Φορ τηις ρεαςον $\bar{\lambda}$ ις οφτεν ςαλλεδ ρυννινγ ςουπλινγ ςονςταντ.
@@
Ιν ςεςτιον 1 ωε πρεςεντεδ τηε ςονςτρυςτιον οφ τηε ιςοςπιν ςψμμετρψ οφ νυςλεαρ φορςες βψ Ηειςενβεργ ανδ Κεμμερ. Ιτ ις βαςεδ ον τηε γρουπ $SU(2)$ ανδ τηε παιρ προτον-νευτρον βελονγς το τηε τωο-διμενςιοναλ φυνδαμενταλ ρεπρεςεντατιον. Τηε διςςοερψ οφ ςτρανγε παρτιςλες βρουγητ τηε νεεδ το ενλαργε τηις ςψμμετρψ ανδ τηε “οβιους” εξτενςιον ωας $SU(2)\rightarrow SU(3)$ βψ ενλαργινγ τηε δουβλετ το α τριπλετ. Ιτ ωας τηε μοδελ προποςεδ βψ Σ. Σαϰατα ιν 1956[77] ιν ωηιςη ηε αδδεδ τηε $\Lambda$-ηψπερον το τηε τωο νυςλεονς το φορμ α φυνδαμενταλ τριπλετ οφ $SU(3)$. Τηις ςςηεμε διδ νοτ φιτ τηε οβςερεδ ηαδρονς ανδ ιν τηε φολλοωινγ ψεαρς μανψ γρουπς ωερε ςονςιδερεδ ας ςανδιδατες φορ τηις λαργερ ςψμμετρψ. Φιναλλψ ιτ τυρνεδ ουτ τηατ Νατυρε ωας μορε ςυβτλε: $SU(3)$ ις τηε ριγητ ανςωερ, βυτ τηε φυνδαμενταλ τριπλετ ις νοτ φορμεδ βψ παρτιςλες ωε ςεε. Ιν 1961 Μ. Γελλ-Μανν ανδ Ψ. Νεέμαν φορμυλατεδ τηε μοδελ ωηιςη βεςαμε ϰνοων ας “Τηε ειγητφολδ ωαψ”[78] ιν ωηιςη τηε λοω-λψινγ ηαδρονς φορμ οςτετς οφ $SU(3)$. Τηε πεςυλιαρ προπερτψ οφ τηις ςςηεμε ις πρεςιςελψ τηε αβςενςε οφ πηψςιςαλ παρτιςλες βελονγινγ το τηε φυνδαμενταλ τριπλετ ρεπρεςεντατιον. Τηις λεδ Γελλ-Μανν ανδ, ινδεπενδεντλψ Γ. Ζωειγ, το προποςε ιν 1964 α ςομποςιτε μοδελ ιν ωηιςη ηαδρονς αρε βουνδ ςτατες οφ ςονςτιτυεντς – ναμεδ “χυαρϰς” βψ Γελλ-Μανν ανδ “αςες” βψ Ζωειγ – ωηιςη φιλλ τηε μιςςινγ τριπλετ[79]. Ωηατ ις λεςς ϰνοων ις τηατ ςιμιλαρ ιδεας ηαδ βεεν εξπρεςςεδ εαρλιερ, Ι βελιεε φιρςτ βψ Λ.Β. Οϰυν[80]. Ηε ςταρτεδ φρομ τηε Σαϰατα μοδελ βυτ ηε ιντερπρετεδ τηε μεμβερς οφ τηε φυνδαμενταλ τριπλετ νοτ ας τηε πηψςιςαλ $p$, $n$ ανδ $\Lambda$, βυτ ας φιςτιτιους παρτιςλες [^45] He used the term “sakatons”.. Ιν φαςτ, ουρ ιδεας ον χυαρϰς φολλοωεδ α ςιμιλαρ τραϑεςτορψ. Ατ φιρςτ τηεψ ωερε ςονςιδερεδ ας ϑυςτ ανοτηερ λαψερ οφ ςομποςιτενεςς ανδ εξπεριμεντς ωερε δεςιγνεδ το διςςοερ τηεμ [^46] There have been even several unsubstantiated claims of discovery.. Ονε οφ τηε μοτιατιονς το βυιλτ ΙΣΡ, τηε φιρςτ ηαδρονις ςολλιδερ, ωας το διςςοερ χυαρϰς. Ηοωεερ, τηε πιςτυρε βεςαμε ςοον μορε ςομπλιςατεδ. Τηε νεεδ το ςατιςφψ τηε ρεχυιρεμεντς οφ Φερμι ςτατιςτιςς, λεδ το τηε ιντροδυςτιον οφ ςολορ[81] ανδ τηε ηψποτηεςις οφ “ςολορ βλινδνεςς” λεδ το τηε νοτιον οφ ςονφινεμεντ. Εεν τοδαψ ωε ηαε νο ρεαλ υνδερςτανδινγ οφ ιτ ανδ ωε ςαννοτ ριγορουςλψ δεριε ιτ φρομ φιρςτ δψναμιςαλ πρινςιπλες [^47] It is one of the unsolved problems endowed with a prize of one million dollars from the Clay Mathematics Institute. For a review of the various ideas around confinement see reference [82].. Ας α ρεςυλτ, ιν ςπιτε οφ ςομε πηενομενολογιςαλ ςυςςεςςες οφ τηε ςιμπλε νον-ρελατιιςτις χυαρϰ μοδελ, μοςτ πεοπλε ςηιφτεδ τηειρ ιεως το ςονςιδερινγ χυαρϰς ας μερε ματηεματιςαλ δειςες το ϰεεπ τραςϰ οφ ςψμμετρψ προπερτιες. Νατυραλλψ, Μ. Γελλ-Μανν, ωηο ςτρονγλψ αδοςατεδ τηις οπινιον, πλαψεδ α μαϑορ ρολε το ιτς γενεραλ ενδορςεμεντ. Τηε ατμοςπηερε ςηανγεδ ωιτη τηε ΣΛΑ῝ ρεςυλτς ον δεεπ ινελαςτις ςςαττερινγ, βυτ Ι ωιλλ διςςυςς τηεμ ιν ςεςτιον 1.
@@
[…]
Στυδψινγ τηε ρενορμαλιςατιον προπερτιες οφ Ψανγ-Μιλλς τηεοριες τυρνεδ ουτ το βε α ερψ ηαρδ ανδ ςομπλεξ προβλεμ, βοτη ςονςεπτυαλλψ ανδ πραςτιςαλλψ. Ον τηε ςονςεπτυαλ ςιδε, νειτηερ τηε ρεςυλτς οφ Φεψνμαν[52] ανδ Φαδδεε ανδ Ποπο[53] νορ τηοςε ον ςποντανεους ςψμμετρψ βρεαϰινγ ωερε ωιδελψ ϰνοων ανδ ῞ελτμαν ηαδ το ρεδιςςοερ εεν τηε βαςις Φεψνμαν ρυλες. Ον τηε πραςτιςαλ ςιδε, τηε νυμβερ οφ τερμς γρεω ερψ φαςτ ανδ ῞ελτμαν ηαδ το δεελοπ α ςομπυτερ προγραμ το ηανδλε τηεμ[120]. Ηε ςαλλεδ ιτ “Σςηοονςςηιπ”,626262“Clean ship” in dutch ανδ ιτ ωας τηε φιρςτ προγραμ οφ ςψμβολις μανιπυλατιονς αππλιεδ το τηεορετιςαλ ηιγη ενεργψ πηψςιςς.